OpenStudy (anonymous):

convert 17/40 to a percent

OpenStudy (anonymous):

17/40 = .425 Move by 2 sides to right 42.5%

OpenStudy (anonymous):

17/40=x/100 solve 1700=40x x=85/2 x=42.5%